# General Hamiltonian description of nonreciprocal interactions

### v3

@@ Appendix D Determining the critical temperature
Appendix E Supplementary results for the Monte-Carlo simulations
This section presents supplementary results for the Monte-Carlo simulations presented in the Main Text. In Sec. E.1, we detail the implementation of the Glauber algorithm under the proposed constraints, derived in Sec. C.2.1. In Sec. E.2, we present numerical results for the individual-update Glauber dynamics that we propose, for Glauber dynamics with selfish energy as proposed in previous works [30, 29, 50], and for Langevin dynamics across various system sizes. Additionally, we examine the outcomes of Glauber dynamics with the Hamiltonian dropping the auxiliary spin (Eq. (104)), and highlight the significant differences from the behavior observed in the constrained Glauber dynamics. Finally, in Sec. E.4, we explore a different vision-cone interaction strength proposed in Ref. [45] and compare the constrained Glauber dynamics with selfish energy dynamics [45] and Langevin dynamics: We find numerically identical behavior obtained from the constrained Glauber dynamics and the Langevin dynamics, whereas the selfish energy approach results in a different behavior.
E.1 Individual-update Glauber algorithm with constraint
In this subsection, we introduce the details of the constrained Glauber algorithm from Sec. C.2.1 that we use in the Main Text. The Hamiltonian is introduced in Eq. (72). We use $\Psi^{(n)}$ to represent the system configuration $\Psi^{(n)}=(\theta_{1}^{(n)},\theta_{2}^{(n)},...\theta_{i}^{(n)},...\theta_{N}^{(n)},\varphi_{1}^{(n)},\varphi_{2}^{(n)},...\varphi_{i}^{(n)},...\varphi_{N}^{(n)})$, where $n$ denotes the iteration step of the Monte-Carlo simulation. The algorithm proceeds as follows:
@@ E.1 Individual-update Glauber algorithm with constraint
    Initial Configuration: Set initial values for the original and auxiliary spins in a uniform configuration, with $\theta_{i}^{(0)}=0$ and $\varphi_{i}^{(0)}=\pi$. This initial configuration satisfies the constraint.
@@
    Single Spin Update: Propose a new configuration by selecting an original spin $\theta_{i}^{(0)}$ randomly and rotating it to a new angle $\theta_{i}^{(0)}\rightarrow\theta_{i}^{(0)}+\delta$, where $\delta$ is drawn from a uniform distribution between $[-\Delta,\Delta]$. As anticipated in Sec. C.2.1, we find empirically that the smaller $\Delta$, the better the agreement with the Langevin steady state.
@@
    Energy Change Calculation: The energy difference between the initial and the proposed configurations is given by $\Delta H=H(\theta_{i}^{(0)}+\delta,\varphi_{i}^{(0)})-H(\theta_{i}^{(0)},\varphi_{i}^{(0)})$. For the XY spins with vision-cone interactions, it is given by
    $\Delta H$ $=$ $-\sum_{j\in\langle ij\rangle}\bigg{\{}J_{ij}(\theta_{i}^{(0)}+\delta)\left[\cos(\theta_{i}^{(0)}+\delta-\theta_{j}^{(0)})+\cos(\varphi_{i}^{(0)}-\theta_{j}^{(0)})\right]$
@@
    $+\sum_{j\in\langle ij\rangle}\bigg{\{}J_{ij}(\theta_{i}^{(0)})\left[\cos(\theta_{i}^{(0)}-\theta_{j}^{(0)})+\cos(\varphi_{i}^{(0)}-\theta_{j}^{(0)})\right]+J_{ji}(\theta_{j}^{(0)})\left[\cos(\theta_{j}^{(0)}-\theta_{i}^{(0)})+\cos(\varphi_{j}^{(0)}-\theta_{i}^{(0)})\right]\bigg{\}}.$
@@
    Transition Probability: Calculate the transition probability $w$ using the individual-update Glauber transition rates at temperature $T$:
    $w=\frac{1}{2}\left(1-\tanh\frac{\Delta H}{2T}\right).$ (103)
@@
    Acceptance/Rejection Step: To accept or reject the proposed new configuration, draw a random number $r$ from a uniform distribution between 00 and $1$. If it is less than the transition probability, i.e., $r<w$, accept the new configuration by updating the selected spin as $\theta_{i}^{(1)}=\theta_{i}^{(0)}+\delta$. Otherwise, reject the update and leave the spin unchanged. Note that all other spins remain unchanged $\theta_{j}^{(1)}=\theta_{j}^{(0)}$, for all $j\neq i$.
@@
    Enforce Constraint: Update the auxiliary spins, $\varphi_{i}^{(1)}=\theta_{i}^{(1)}+\pi$ to enforce the constraint. We thus obtain the next-step configuration $\Psi^{(1)}$.
  7. 7.
    Repeat: Iterate the procedure for a large number of steps. In our numerical simulations, we perform $N^{5}$ iterations where $N$ is the system size.
We note in passing that a similar algorithm can be constructed for the joint-update scheme in a straightforward way, following the analysis in Sec. C.2.2.
Figure 7: Average magnetization $\langle m\rangle$ (a,b,c) and specific heat $C_{v}$ (d,e,f) obtained from numerical simulations of Langevin dynamics, individual-update constrained Glauber dynamics, and selfish-energy Glauber dynamics as functions of temperature $T$. While the error bars of the magnetization decrease with the linear system size $L$ for the selfish-energy Glauber dynamics, they do not show a clear system-size dependence near the critical point for constrained Glauber dynamics and Langevin dynamics. The remaining parameter values are the same as in Fig. 3.
E.2 Effects of system size
In this subsection, we investigate the effects of system size as shown in Fig. 7. The average magnetization and specific heat from both Langevin dynamics and the individual-update constrained Glauber dynamics that we propose here exhibit similar behavior across ordered and disordered phases at different system sizes, differing significantly from the results of selfish-energy Glauber dynamics proposed in previous works [30, 29, 50].
For the selfish-energy Glauber dynamics, the error bars of the magnetization decrease as the system size $L$ increases. A similar trend is observed for constrained Glauber dynamics and Langevin dynamics at temperatures away from the critical point. However, near the critical point, no clear correlation between the error bars and the system size is observed.
@@ E.2 Effects of system size
E.3 Glauber dynamics without auxiliary spins
@@
Figure 9: Average magnetization $\langle m\rangle$ in the steady state on a square lattice of linear dimension $L=100$, as a function of temperature $T$ for individual-update constrained Glauber (blue dots), selfish-energy Glauber (red dots), and Langevin (green squares) dynamics. The results indicate that in the long-time steady-state, the constrained Glauber dynamics effectively simulates the behavior of the original Langevin dynamics; the agreement is particularly good away from the critical temperature $T_{c}=0.54J$. Close to the critical point, fluctuations prevent reaching a steady state within the $6\times 10^{6}$ ($20\times L^{4}$) iterations used for Langevin (Constrained Glauber), but the critical temperatures agree within the statistical error bars $\Delta T_{c}=0.02J$ (b). By contrast, selfish-energy Glauber dynamics results in significant differences compared to Langevin dynamics, both in the critical region and away from it (a). We use the vision-cone interaction strength in Eq. (105) with $\sigma=1$, along with the semi-width $\Delta=0.5$ for constrained Glauber dynamics, and a time step of $\delta t=0.1$ for the Langevin dynamics.
E.4 Comparison between selfish energy and constrained Glauber dynamics for an alternative vision-cone interaction
@@
$E_{i}=-\sum_{j\in\langle ij\rangle}\exp{[\sigma\cos{(\theta_{i}-\psi_{ij})}]}\cos(\theta_{i}-\theta_{j}).$ ) . (106)
@@
$\dot{\theta_{i}}=-\sum_{j\in\langle ij\rangle}\exp{[\sigma\cos{(\theta_{i}-\psi_{ij})}]}\sin(\theta_{i}-\theta_{j})+\sqrt{2T}\eta_{i}(t),$
$\dot{\varphi_{i}}=\sum_{j\in\langle ij\rangle}\exp{[\sigma\cos{(\theta_{i}-\psi_{ij})}]}\sin(\varphi_{i}-\varphi_{j})+\sqrt{2T}\eta_{i}(t).$ ) . (107)
@@
The average steady-state magnetization for the three alternative approaches at system size $L=100$ is shown in Fig. 9, for $\sigma=1$. For the constrained Glauber dynamics, we use semi-width $\Delta=0.5$ and $20\times L^{4}$ iterations; A time step $\delta t=0.1$ is used for Langevin dynamics.

### v4

@@ Appendix E Determining the critical temperature for the vision-cone model
Appendix F Supplementary results for the Monte-Carlo simulations
This section presents supplementary results for the Monte-Carlo simulations presented in the Main Text. In Sec. F.1, we detail the implementation of the Glauber algorithm under the proposed constraints, derived in Sec. F.4. In Sec. F.3, we present numerical results for the individual-update Glauber dynamics that we propose, for Glauber dynamics with selfish energy as proposed in previous works [5, 51, 15], and for Langevin dynamics across various system sizes. Additionally, we examine the outcomes of Glauber dynamics with the Hamiltonian dropping the auxiliary spin (Eq. (196)), and highlight the significant differences from the behavior observed in the constrained Glauber dynamics. Finally, in Sec. F.5, we explore a different vision-cone interaction strength proposed in Ref. [67] and compare the constrained Glauber dynamics with selfish energy dynamics [67] and Langevin dynamics: We find numerically identical behavior obtained from the constrained Glauber dynamics and the Langevin dynamics, whereas the selfish energy approach results in a different behavior.
F.1 Constrained Glauber algorithm
In this subsection, we introduce the details of the constrained Glauber algorithm from Sec. D.2 that we use in the Main Text. The Hamiltonian is introduced in Eq. (141). We use $\Psi^{(n)}$ to represent the system configuration $\Psi^{(n)}=(\theta_{1}^{(n)},\theta_{2}^{(n)},...\theta_{i}^{(n)},...\theta_{N}^{(n)},\varphi_{1}^{(n)},\varphi_{2}^{(n)},...\varphi_{i}^{(n)},...\varphi_{N}^{(n)})$, where $n$ denotes the iteration step of the Monte-Carlo simulation. The algorithm proceeds as follows:
@@ F.1 Constrained Glauber algorithm
    Initial Configuration: Set initial values for the original and auxiliary spins in a uniform configuration, with $\theta_{i}^{(0)}=0$ and $\varphi_{i}^{(0)}=\pi$. This initial configuration satisfies the embedding constraint.
@@
    Original Spin Update: Propose a new configuration by selecting an original spin uniformly at random and rotating it to a new angle $\theta_{i}^{(0)}\rightarrow\theta_{i}^{(0)}+\delta$, where $\delta$ is drawn from a uniform distribution between $\Delta,\Delta]$. As anticipated in Sec. D.2, we find that the smaller $\Delta$, the better the agreement with the Langevin steady state. The energy of the new configuration reads as $H(\theta_{i}^{(0)}+\delta,\varphi^{(0)}_{i})$ where $H$ is the embedding Hamiltonian.
@@
    Auxiliary Spin Update: We also update the the auxiliary spin $\varphi_{i}$ by the same angle $\delta$ to ensure that the constraint is preserved at each step, i.e., $\varphi_{i}^{(0)}\rightarrow\varphi_{i}^{(0)}+\delta$. The energy of this configuration is $H(\theta_{i}^{(0)},\varphi_{i}^{(0)}+\delta)$.
  4. 4.
    Energy Change Calculation: The energy difference between the initial and the proposed configurations is given by $\Delta E=\frac{1}{2}[H(\theta_{i}^{(0)}+\delta,\varphi_{i}^{(0)})-H(\theta_{i}^{(0)},\varphi_{i}^{(0)}+\delta)]$. For the XY spins with vision-cone interactions, it is given by
    $\Delta E$ $=$ $-\frac{1}{2}\sum_{j\in\langle ij\rangle}\bigg\{J_{ij}(\theta_{i}^{(0)}+\delta)\left[\cos(\theta_{i}^{(0)}+\delta-\theta_{j}^{(0)})+\cos(\varphi_{i}^{(0)}-\theta_{j}^{(0)})\right]$
@@
    $+\frac{1}{2}\sum_{j\in\langle ij\rangle}\bigg\{J_{ij}(\theta_{i}^{(0)})\left[\cos(\theta_{i}^{(0)}-\theta_{j}^{(0)})+\cos(\varphi_{i}^{(0)}+\delta-\theta_{j}^{(0)})\right]+J_{ji}(\theta_{j}^{(0)})\left[\cos(\theta_{j}^{(0)}-\theta_{i}^{(0)})+\cos(\varphi_{j}^{(0)}-\theta_{i}^{(0)})\right]\bigg\}.$
  5. 5.
    Transition Probability: Calculate the transition probability $w$ using the Glauber transition rates at temperature $T$:
    $w=\frac{1}{2}\left(1-\tanh\frac{\Delta E}{2T}\right).$ (193)
  6. 6.
    Acceptance/Rejection Step: To accept or reject the proposed update, draw a random number $r$ from a uniform distribution between 0 and $1$. If it is less than the transition probability, i.e., $r<w$, accept the new configuration by updating the spins on the selected site as $\theta_{i}^{(1)}=\theta_{i}^{(0)}+\delta$, and $\varphi_{i}^{(1)}=\varphi_{i}^{(0)}+\delta$. Otherwise, reject the update and leave the spin unchanged. Note that all other spins remain unchanged $\theta_{j}^{(1)}=\theta_{j}^{(0)}$, for all $j\neq i$.
  7. 7.
    Repeat: Iterate the procedure for a large number of steps. In our numerical simulations, we perform $200L^{4}$ iterations where $L$ is the linear dimension.
F.2 Constrained Glauber algorithm after eliminating the auxiliary spins
In Sec. D.2.3, we have shown that the auxiliary spins can be eliminated from the expression for the transition rate, and hence we can simulate the constrained Glauber dynamics using only the original spins $\theta_{i}$. The system configuration is now $\Psi^{(n)}=(\theta_{1}^{(n)},\theta_{2},...\theta_{N}^{(n)})$.
The algorithm proceeds as follows:
  1. 1.
    Initial Configuration: Set initial values for the original spins in a uniform configuration, with $\theta_{i}^{(0)}=0$. This initial configuration satisfies the constraint.
  2. 2.
    Spin Update: Propose a new configuration by selecting a spin uniformly at random and rotating it to a new angle $\theta_{i}^{(0)}\rightarrow\theta_{i}^{(0)}+\delta$, where $\delta$ is drawn from a uniform distribution between $\Delta,\Delta]$.
  3. 3.
    Energy Change Calculation: The energy difference between the initial and the proposed configurations is given by Eq. (189), which reads as
    $\Delta E=-\frac{1}{2}\sum_{j\in\langle ij\rangle}\left[J_{ij}(\theta_{i}^{(0)}+\delta)+J_{ij}(\theta_{i}^{(0)})\right]\,\left[\cos(\theta_{i}^{(0)}+\delta-\theta_{j}^{(0)})-\cos(\theta_{i}^{(0)}-\theta_{j}^{(0)})\right].$ (194)
@@ F.2 Constrained Glauber algorithm after eliminating the auxiliary spins
    Transition Probability: Calculate the transition probability $w$ using the Glauber transition rates at temperature $T$:
    $w=\frac{1}{2}\left(1-\tanh\frac{\Delta E}{2T}\right).$ (195)
@@
    Acceptance/Rejection Step: To accept or reject the proposed new configuration, draw a random number $r$ from a uniform distribution between 0 and $1$. If it is less than the transition probability, i.e., $r<w$, accept the new configuration by updating the selected spin as $\theta_{i}^{(1)}=\theta_{i}^{(0)}+\delta$. Otherwise, reject the update and leave the spin unchanged. Note that all other spins remain unchanged $\theta_{j}^{(1)}=\theta_{j}^{(0)}$, for all $j\neq i$.
@@
    Repeat: Iterate the procedure for a large number of steps. In our numerical simulations, we perform $200L^{4}$ iterations where $L$ is the linear dimension.
Figure 7: Average magnetization $\langle m\rangle$ (a,b,c), specific heat $C_{v}$ (d,e,f), and error bars of average magnetization $\sigma_{m}$ (g,h,i) obtained from numerical simulations of Langevin dynamics, constrained Glauber dynamics, and selfish-energy Glauber dynamics as functions of temperature $T$. Away from the transition, the magnetization fluctuations are self-averaging, so the standard deviation (error bars) decreases with increasing linear size $L$ for all three dynamics. Close to the critical point, the peak of error bars increases with increasing $L$ for both Langevin and constrained Glauber dynamics, whereas the selfish-energy Glauber dynamics shows decreasing error bars upon increasing $L$ (at least up to $L=100$). The results indicate that the selfish-energy Glauber dynamics samples a different stationary distribution rather than reproducing the critical behavior of the Langevin dynamics. The remaining parameter values are the same as in Fig. 3.
F.3 Finite-size effects
In this subsection, we investigate the effects of system size as shown in Fig. 7. The average magnetization and specific heat from both Langevin dynamics and the constrained Glauber dynamics that we propose here exhibit similar behavior across ordered and disordered phases at different system sizes, differing significantly from the results of selfish-energy Glauber dynamics proposed in previous works [5, 51, 15].
To better characterize the fluctuations in the magnetization, we also examine the standard deviation $\sigma_{m}=\sqrt{\langle m^{2}\rangle-\langle m\rangle^{2}}$ across different system sizes (shown as error bars in the panels of the first row in Fig. 7). Both Langevin and constrained Glauber dynamics exhibit similar scaling behavior, which differs markedly from that of the selfish-energy Glauber dynamics. Away from the critical temperature, the error bars decrease with increasing system size $L$ in all three methods, as expected due to self-averaging in large systems. However, near the critical temperature, Langevin and constrained Glauber dynamics show increasing error bars with system size, reflecting critical fluctuations that grow in the thermodynamic limit. By contrast, the selfish-energy Glauber dynamics exhibits the opposite trend, with decreasing error bars near the transition, suggesting a different critical behavior. This discrepancy further supports the usefulness of our constrained Hamiltonian embedding approach in capturing the correct nonequilibrium physics of nonreciprocal systems.
@@ F.3 Finite-size effects
F.4 Glauber dynamics without auxiliary spins
@@
Figure 9: Average magnetization $\langle m\rangle$ in the steady state on a square lattice of linear dimension $L=100$, as a function of temperature $T$ for constrained Glauber (blue dots), selfish-energy Glauber (green dots), and Langevin (red squares) dynamics. The results show that in the long-time steady state, the constrained Glauber dynamics accurately reproduces the behavior of the Langevin dynamics and captures the same critical temperature, $T_{c}/J\approx 0.74$ (a), whereas the selfish-energy Glauber dynamics predicts a significantly different critical temperature, $T_{c}/J\approx 1.07$ (b). We use the vision-cone interaction strength in Eq. (197) with $\sigma=1$, the semi-width $\Delta=0.2$ for constrained Glauber dynamics, and a time step of $\delta t=0.01$ for Langevin dynamics. We perform $200L^{4}$ iterations ($100L^{2}$ time steps) for constrained Glauber (Langevin) dynamics and $70L^{4}$ iterations for selfish-energy Glauber dynamics. The ensemble consists of $\mathcal{N}=100$ independent trajectories.
F.5 Soft-vision-cone interactions: Comparison between selfish energy and constrained Glauber dynamics
@@
[…]
@@
[…]
@@
The average steady-state magnetization for the three alternative approaches at system size $L=100$ is shown in Fig. 9, for $\sigma=1$. For the constrained Glauber dynamics, we use a semi-width $\Delta=0.2$ and $200\times L^{4}$ iterations; a time step $\delta t=0.01$ is used for Langevin dynamics.